Two athletes are competing in a race. One averages 10mph, and the other 12 mph. If the faster runner finishes one minute sooner than the other how long is the race?
I am a parent trying to help my 13 year old with her algebra homework.

Answer by stanbon(75887)

Two athletes are competing in a race. One averages 10mph, and the other 12 mph. If the faster runner finishes one minute sooner than the other how long is the race?
---------
Slower Runner DATA:
rate = 10 mph ; distance = x miles ; time = distance/rate = x/10 hrs
-------
Faster Runner DATA:
rate = 12 mph ; distance = x miles ; time = dist/rate = x/12 hrs
------
Note:
x/10 hrs = 60(x/10) = 6x minutes
x/12 hrs = 60(x/12) = 5x minutes
---------------------
Equation:
slower runner time - faster runner time = 1 minute
6x - 5x = 1 min
x = 1 mile
Ans: The length of the race is 1 mile
=====
Thanks for helping your daughter.
Cheers,
Stan H.